ISS On-Orbit Status 25 June 2004

All ISS systems continue to function nominally except those noted previously or below.

Last night's Orlan EVA was aborted shortly after the crew's egress from the DC1 airlock. [After on-time hatch opening at 5:56pm EDT, Russian flight controllers at TsUP/Moscow called the spacewalk off, ordering Gennady Padalka (EV1) and Mike Fincke (EV2) back inside, when they noticed (and the crew confirmed) a larger-than-expected pressure drop in the primary oxygen tank of Fincke's Orlan-M spacesuit. After ingress, the crewmembers hooked up to suit umbilicals and closed the hatch. After preliminary checks, TsUP announced a "No Go" for the EVA at ~6:59pm. Total duration of the excursion was 14 min 22 sec. Hatches between Russian (RS) and U.S. (USOS) segments were open at 8:27pm. The crew then reversed all pre-EVA preparation steps in RS and USOS.]

Further troubleshooting and testing on Mike Fincke's primary O2 supply indicates that the suit's injector mode selector valve did not seat fully when closed, allowing O2 to flow rapidly thru the suit unnoticed by the crewmember. [The injector activation LED (light-emitting diode) indicator on Mike's helmet did not illuminate during the injector flow when the O2 selector handle was not moved hard enough, as was a subsequent test confirmed. The suit pressure did not increase during this period because the opening of the pneumatically controlled relief valve is linked to injector activation, venting O2 at the same rate as the injector supplies. Therefore, during the airlock depressurization Mike and the ground did not notice the continuing high rate of O2 flowing through the suit and the resulting tank pressure bleed-down until immediately after hatch opening.]

The new (preliminary) EVA date (pending thermal analyses) is Wednesday, 6/30, with hatch opening at ~5:40pm EDT and hatch closing at ~11:40pm. There is general agreement that Mike Fincke's suit #26 (blue markings) continues to be Go, when equipped with a fresh O2 bottle.

Wake-up time for the crew on today's shifted cycle was 8:00am EDT, with sleep time set for tonight at the regular 5:30pm. For Saturday and Sunday, the crew thus returns to the regular 2:00am-to-5:30pm cycle.

Shortly after post-sleep activities, CDR Padalka and FE/SO Fincke completed their post-EVA session of the Russian crew health-monitoring program's medical assessment MO-9/Biochemical Urinalysis. Gennady then stowed the Urolux hardware. [MO-9 is conducted regularly every 30 days (and also before and after EVAs) and is one of five nominal Russian medical tests adopted by NASA for US crewmembers for IMG PHS (Integrated Medical Group/Periodic Health Status) evaluation as part of the "PHS/Without Blood Labs" exam. The analysis uses the sophisticated in-vitro diagnostic apparatus "Urolux" developed originally for the Mir program. Afterwards, the data are entered in the medical equipment computer (MEC)'s special IFEP software (In-Flight Examination Program).]

At 10:40am, Padalka repressurized the station's cabin atmosphere with 15 Torr (mmHg) of fresh air from Progress 14P, from the second tank (SrPK).

The CDR reactivated the Sputnik-SM ham radio hardware in the Service Module (SM).

Afterwards, Gennady also collected the radiation measurements taken by the "Pille-MKS" dosimeter sensors in the two Orlan suits and the static background monitor in the SM.

FE/SO Fincke took the periodic CO2 partial pressure readings in the SM and Lab using the U.S. CDMK (CO2 monitor kit), for calldown to MCC-Houston (along with the battery status) for use in trending analyses. The ground also activated the MCA (Major Constituent Analyzer) for the weekly readings.

Mike completed the regular daily maintenance/inspection of the SM's SOZh environment control & life support systems, including routine ASU toilet system replacements.

The crew discussed specifics of the crew's Orlan-M suit operations with ground specialists in a two-hour tagup via S-band at 11:30am

The crew's regular weekly teleconference with the Flight Director is scheduled for 3:40pm (S-band/audio).

Today's CEO (Crew Earth Observations) photo targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were Los Angeles, California (ISS approached from the NW with the coast near nadir. Shooting as many images as needed to capture the margin of this large urban region), and von Karman cloud vortices, Baja California (ISS passed down the inland side of the Gulf of California. Looking right to the islands beyond the Baja peninsula for possible cloud eddies. Conditions were setting up for these to form).

CEO images can be viewed at the websites http://eol.jsc.nasa.gov and http://earthobservatory.nasa.gov.

See also the website "Space Station Challenge" at http://voyager.cet.edu/iss/.

U.S. & Russian Segment Status (as of today, 12:30pm EDT)

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

Elektron O2 generator is On. Vozdukh CO2 scrubber is Off. U.S. CDRA CO2 scrubber is Off. TCCS (trace contaminant control subsystem) is operating. SM Gas Analyzer has been calibrated and is used for ppO2 & ppCO2 monitoring. MCA (major constituents analyzer) is in Life Extending Mode (LEM). BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode. RS air conditioner SKV-1 is Off, SKV-2 is Off (SM panel mods completed 4/8; SKV-2 activation failed 4/20; is now functioning again). SFOG slot #2 fan suspect (not usable).


SM Working Compartment: Pressure (mmHg) -- 742; temperature (deg C) -- 25.2; ppO2 (mmHg) -- 153.0; ppCO2 (mmHg) -- 2.5.
SM Transfer Compartment: Pressure (mmHg) -- 741; temperature (deg C) -- 19.8.
FGB Cabin: Pressure (mmHg) -- 740; temperature (deg C) -- 21.7.
Node: Pressure (mmHg) -- 740.71; temperature (deg C) -- 23.6 (shell); ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
U.S. Lab: Pressure (mmHg) -- 742.59; temperature (deg C) -- 24.9; ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
Joint Airlock (Equip. Lock): Pressure (mmHg) -- 742.69; temperature (deg C) -- 22.8; shell heater temp (deg C) -- n/a, ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.


(n/a = data not available)


PMA-1: Shell heater temp (deg C) -- n/a
PMA-2: Shell heater temp (deg C) -- n/a.

Electrical Power Systems (EPS): Both P6 channels fully operational. BGA (beta gimbal assembly) 2B and 4B both in directed position (Blind mode, non solar-tracking, biased for drag reduction). SM batteries: All batteries (8) are on line in "Partial Charge" mode. FGB batteries: Battery #5 is off line; all other batteries (5) are on line in "Partial Charge" mode.

Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.

Command & Data Handling Systems:

C&C-1 MDM is prime, C&C-2 is backup, and C&C-3 is in standby.
GNC-2 MDM is prime; GNC-1 is backup.
INT-2 is operating; INT-1 is Off.
EXT-1 is On (primary), EXT-2 is Off (backup).
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-1 MDM is Off; PL-2 MDM is Operational.
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22/03).
SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22/03).
FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.

Propulsion System:

Total propellant load available: 3963 kg (8737 lb) as of 6/2410/04; [SM(552) + FGB(2772) + Progress M(639)]. (Capabilities: SM -- 860 kg; FGB -- 6120 kg).

Attitude Control Systems:

2 CMGs on-line (CMG-1 failed, since 6/6/02; CMG-2's RPC-17 failed 4/21/04).
State vector source -- U.S. SIGI-1 (GPS)
Attitude source -- U.S. SIGI-1 (GPS)
Angular rate source -- RGA-1

Flight Attitude:

LVLH XVV (local vertical/local horizontal = "earth-fixed": z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -9 deg, roll: 0 deg]), with CMG/TA (thruster assist) Momentum Management, until 6/28, following the EVA.

Communications & Tracking Systems:

FGB MDM-1 is powered Off; FGB MDM-2 is operational.
All other Russian communications & tracking systems are nominal.
S-band is operating nominally (on string 2).
Ku-band is operating nominally (may require a mask).
Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
Video subsystem operating nominally.
HCOR (high-rate communications outage recorder) is operating nominally.

Robotics:

SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
MBS: KA (keep alive) power on both strings.
MT: latched and mated at WS4.
POA: KA power on both strings.
RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit (as of this morning, 9:33am EDT [= epoch]):

Mean altitude -- 360.8 km
Apogee -- 364.5 km
Perigee -- 357.1 km
Period -- 91.8 min.
Inclination (to Equator) -- 51.6327 deg
Eccentricity -- 0.0005462
Solar Beta Angle -- 6.8 deg
Orbits per 24-hr. day -- 15.69
Mean altitude loss in last 24 hours -- 70 m
Revolutions since FGB/Zarya launch (Nov. '98) -- 31963

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html.

Source: NASA